OpenStudy (anonymous):

Perimeter=70in. The sides are 21, 20 and x What is x?

OpenStudy (anonymous):

perimeter of what?

OpenStudy (anonymous):

A triangle

OpenStudy (anonymous):

70-21-20=29

OpenStudy (anonymous):

ooh ok \[20+21+x=70\\ 41+x=70\] solve for \(x\) in one step